OpenStudy (anonymous):

What sine function represents an amplitude of 1, a period of 2π, a horizontal shift of π, and a vertical shift of −4?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you have one that looks like \[\sin(x-\pi)-4\]?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

the \(-\pi\) insides the arguement of the function will shift it to the RIGHT \(\pi\) units and the \(-4\) out at the end shifts it down 4

OpenStudy (misty1212):

this really has nothing to do with trig it is the same for any function \(f\)

OpenStudy (anonymous):

thank u!